OpenStudy (anonymous):

what is sin (-13pi/4) what is sin (-13pi/4) @Mathematics

OpenStudy (anonymous):

tyler can you help me with these two pls cos ( cos (5pi/3) cos(-270 degrees) and please show work

OpenStudy (anonymous):

I can help you with the first one.

OpenStudy (anonymous):

First, change \[- \frac{13 \pi}{4} \] into degrees to make things easy.

OpenStudy (anonymous):

Do you know how to do that?

OpenStudy (anonymous):

is it -13/4 times 180

OpenStudy (anonymous):

Almost

OpenStudy (anonymous):

\[- \frac{13 \pi}{4} * \frac{180}{\pi}\]

OpenStudy (anonymous):

so we get -585 degrees?

OpenStudy (anonymous):

So essentially yes because the pi's cancel.

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

ok what do we do with the degree

OpenStudy (anonymous):

-585 is more than one revolution. So we can add a revolution (or 360 degrees) to -585

OpenStudy (anonymous):

-585 + 360 = ?

OpenStudy (anonymous):

-225

OpenStudy (anonymous):

Good.

OpenStudy (anonymous):

Now i like to deal with positive degrees. So to turn -225 into positive degrees. You do this \[-225 + 360\] Which is what?

OpenStudy (anonymous):

135

OpenStudy (anonymous):

Ok so saying \[\sin(- \frac{13\pi}{4})\] is like saying \[\sin(135)\] which is much easier to work with.

OpenStudy (anonymous):

so thats the final answer, 135?

OpenStudy (anonymous):

Well what is sin(135) ?

OpenStudy (anonymous):

sqrt2/2

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

awesome, why cant you help me with the other two...you explain better than my teacher lol

OpenStudy (anonymous):

I can.. i wanted to start with the first one lol

OpenStudy (anonymous):

ok kool

OpenStudy (anonymous):

so for cos(5pi/3) we convert to degrees first right

OpenStudy (anonymous):

ok so lets do \[\cos (\frac{5\pi}{3})\]

OpenStudy (anonymous):

and get 300 degrees

OpenStudy (anonymous):

Ok good.

OpenStudy (anonymous):

do i do any subtraction...not sure

OpenStudy (anonymous):

Well, you can.

OpenStudy (anonymous):

so do i use 180 or 360

OpenStudy (anonymous):

You should know what sin cos and tan of pi/3, pi/4 and pi/6 are.

OpenStudy (anonymous):

So 300 is in the 4 th quadrant and you want a angle in the first so subtract 300 from 360

OpenStudy (anonymous):

So \[\cos(300) = \cos(60)\]

OpenStudy (anonymous):

and cos(60) is 1/2?

OpenStudy (anonymous):

Now what is cos (60)? or pi/3

OpenStudy (anonymous):

1/2?

OpenStudy (anonymous):

Correct its 1/2

OpenStudy (anonymous):

good stuff man

OpenStudy (anonymous):

cos(-270) we add 360 and get 90 and cos 90 is 0?

OpenStudy (anonymous):

now \[\cos(-270)\]

OpenStudy (anonymous):

You got it!! You got the hang of it now!! :)

OpenStudy (anonymous):

yeah thanks very much for the help

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

ok i have another one Find sin of angle if sin of angle < 0 and cos of angle = 2/3

OpenStudy (anonymous):

1 and 4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

i see how the amplitude is one

OpenStudy (anonymous):

how do i find the period